OpenStudy (anonymous):

i need help im really bad at math and i still cant figure out slope intercept form can you help me on this problem: i have to write an equation in slope intercept form that passes through the given points. 2,-2 , (5,7)

OpenStudy (anonymous):

i really need help

OpenStudy (anonymous):

http://www.purplemath.com/modules/strtlneq.htm

OpenStudy (anonymous):

this explains thoroughly what is slope intercept for is.

OpenStudy (anonymous):

i've already been there but i didnt understand it that well thats how bad at math i am

OpenStudy (anonymous):

Try to read everything not just skim through it. It actually explains very good.

OpenStudy (anonymous):

Slope intercept form is: y=mx+b m represents the slope of the line, and b is the y-intercept. When you're given two points and want to find the slope between them, you use the expression Y2-Y1 7-(-2) 9 ------ ------- = ---- = 3 X2-X1 in this case, you would put 5-2 3 Once you've solved this expression, you have the slope of the line, or m. The slope of this line is 3. So now, you need to find b, and you can do that by using one of the coordinate pairs and plug it into the equation and solve for b. 7=3(5)+b Now you know that b=-8, so your slope intercept equation 7=15+b would be y=3X-8. I hope that helped! 7-15=15-15+b -8=b

OpenStudy (anonymous):

thank you so much for your really really good answer

OpenStudy (anonymous):

Glad to help!